Madam Ling was 28 years old when her son was born. If the product of their ages now is 480,how old is her son now?
Answer by ankor@dixie-net.com(22740)

Madam Ling was 28 years old when her son was born.
If the product of their ages now is 480, how old is her son now?
:
Let s = son's age now
we know
s + 28 = Ling's age now
then the product of their ages:
s(s+28) = 480
s^2 + 28s - 480 = 0
Factors to
(s+40)(s-12) = 0
the positive solution
s = 12 yrs old is the son's present age
:
:
Check this yourself, find Ling's age and multiply by 12